how is 10 divided by 2 =2??????????

Grade:10

1 Answers

Deep Test tutors
9 years ago
2/10 

= TWO/TEN 

=WO/EN 

W 23rd letter , o 15 th letter , E- 5th ,N-14 th 

=(23+15 ) / (15+14) 

=38/19 

=2